OpenStudy (jango_in_dtown):

A number when divided by 97 leaves remainder 73. Then same number when divided by 83 has quotient 42. Find the remainder in the second case.

OpenStudy (jango_in_dtown):

@ganeshie8

OpenStudy (jango_in_dtown):

@mathmate

OpenStudy (jango_in_dtown):

@Loser66

OpenStudy (jango_in_dtown):

@confluxepic

OpenStudy (jango_in_dtown):

@pooja195

OpenStudy (jango_in_dtown):

@Kainui

OpenStudy (jango_in_dtown):

@imqwerty

OpenStudy (jango_in_dtown):

@loser66 hi

OpenStudy (loser66):

Hi, I didn't figure out how to solve it yet!! hehehe my brain is useless.

OpenStudy (loser66):

I have a bunch of solutions so that I don't know how to narrow down the answer.

OpenStudy (jango_in_dtown):

options are 47,69,79,53

OpenStudy (loser66):

oh, so 79 works

OpenStudy (loser66):

If you have options, that is easy. just put it into the equation to get the answer

OpenStudy (loser66):

Like this: Let n is the given number, then \(n = 97* t + 73\\n=83*s+ r\) So, you have \(97t+73 = 83s+r\) then \(97t-83s+73=r\) We know that \(0\leq r\leq 82\), so we narrow down \(97t-83s =r -73\leq 9\)

OpenStudy (jango_in_dtown):

in the second case, s=42

OpenStudy (loser66):

Now, start from 9 97t -83s =9, and (97,83)=1 and then \(97 = (1)83 +14\rightarrow 14= 97-83\\83=(5)14+13\rightarrow 13=83-(5)14\\14=13+1\rightarrow 1=14-13\) Go backwark to get \(1=14-(83-(5)14=(6)14-83\\1=(6)(97-83)-83=(6)97-(7)83\)

OpenStudy (loser66):

multiple both sides by 9, you have \(9=(54)97-(63)83\) Hence that gives you t = 54, s = 63 and r = 82

OpenStudy (loser66):

Take down 1 unit, that is multiple both sides by 8, you have \(8= (48)97-(56)83\), that gives you t=48, s=56 and r = 81

OpenStudy (loser66):

if you down 2 more units, you have \(6= (36)97-(42)83\) so t =36, s =42, r = 79 (one of your options) Now test n = 36* 97 +73= 3565 n = 42*83 +79= 3565 Bingo

OpenStudy (jango_in_dtown):

hA ha ha ... great... :)

OpenStudy (loser66):

hehehe... not great at all. It took me so much time to work!!

imqwerty (imqwerty):

n/97 =97q+73 n/83=83(42)+r n/83 =97(35)+91+r Now 91+r when divided by 97 must give remainder 73 and also r must be less than 83 which gives r=79